# Leave-One-Out Neighborhood Smoothing for Graphons: Berry-Esseen Bounds, Confidence Intervals, and Honest Tuning

### v1

@@ Abstract
Neighborhood smoothing methods achieve minimax-optimal rates for estimating edge probabilities under graphon models, but their use for statistical inference has remained limited. The main obstacle is that classical neighborhood smoothers select data-driven neighborhoods and average edges using the same adjacency matrix, inducing complex dependencies that invalidate standard concentration and normal approximation arguments.
We introduce a leave-one-out modification of neighborhood smoothing for undirected simple graphs. When estimating a single entry $P_{ij}$, the neighborhood of node $i$ is constructed from an adjacency matrix in which the $j$th row and column are set to zero, thereby decoupling neighborhood selection from the edges being averaged. We show that this construction restores conditional independence of the centered summands, enabling the use of classical probabilistic tools for inference.
Under piecewise Lipschitz graphon assumptions and logarithmic degree growth, we derive variance-adaptive concentration inequalities based on Bousquet’s inequality and establish Berry-Esseen bounds with explicit rates for the normalized estimation error. These results yield both finite-sample and asymptotic confidence intervals for individual edge probabilities. The same leave-one-out structure also supports an honest cross-validation scheme for tuning parameter selection, for which we prove an oracle inequality. The proposed estimator retains the optimal row-wise mean-squared error rates of classical neighborhood smoothing while providing valid entrywise uncertainty quantification.
@@ 1 Introduction
Graphon models provide a canonical framework for nonparametric analysis of network data, representing latent edge probabilities through a measurable kernel on the unit square (Borgs et al., 2008; Lovász, 2012). Over the past decade, a substantial literature has established sharp minimax rates for estimating the probability matrix $P$ or the underlying graphon $h$ from a single observed adjacency matrix, under a variety of smoothness and structural assumptions (Bickel et al., 2011; Wolfe et al., 2013; Gao et al., 2015; Chan et al., 2014). From the perspective of point estimation, the theory is now mature.
What remains largely unresolved is statistical inference for individual edge probabilities. Given an estimator $\widehat{P}_{ij}$, it is unclear how to attach a valid measure of uncertainty-such as a confidence interval-to this estimate under nonparametric graphon models. Existing inferential theory for networks has focused almost exclusively on low-dimensional or global functionals, including eigenvalues, eigenvectors, and spectral embeddings (Lei and Rinaldo, 2015; Athreya et al., 2018; Tang et al., 2017; Capitaine et al., 2019). Entrywise inference for $P_{ij}$, despite its natural interpretation and relevance for downstream tasks, has received comparatively little attention.
This gap is not merely technical. It is structural.
Neighborhood smoothing estimators, such as those of Zhang et al. (2017), achieve optimal mean-squared error by averaging observed edges over data-driven neighborhoods of similar nodes. However, these procedures reuse the adjacency matrix twice: first to construct neighborhoods, and then to average edges within those neighborhoods. As a consequence, the random neighborhood weights depend on the very edges being averaged. Even after conditioning on the underlying probability matrix, the centered summands in the estimator are no longer independent. This dependence invalidates the assumptions underlying classical concentration inequalities and central limit theorems, and it is the fundamental reason why standard inferential tools fail for neighborhood smoothers.
This paper shows that this obstruction can be removed by a simple but principled modification.
We introduce leave-one-out (LOO) neighborhood smoothing. When estimating a single entry $P_{ij}$, we construct the neighborhood of node $i$ using a modified adjacency matrix in which the $j$th row and column are set to zero. Distances and neighborhood selection are computed using this modified matrix, after which the estimator averages the original edges $\{A_{kj}:k\in\mathcal{N}_{i}^{(-j)}\}$. By construction, conditional on the probability matrix, the random neighborhood $\mathcal{N}_{i}^{(-j)}$ is independent of the edges used to form the estimate.
This decoupling is the central insight of the paper. It restores conditional independence of the centered summands and makes classical probabilistic tools available for network inference. In particular, we derive variance-adaptive concentration inequalities based on Bousquet’s inequality (Bousquet, 2002) and establish Berry-Esseen bounds with explicit rates for the normalized estimation error (Petrov, 1975; Vershynin, 2018). These results yield both finite-sample and asymptotic confidence intervals for individual entries of the probability matrix.
The same leave-one-out structure also enables honest tuning via cross-validation. Because the LOO predictor for $P_{ij}$ is independent of the held-out edge $A_{ij}$ conditional on the constructed neighborhoods, squared prediction error provides an unbiased estimate of prediction risk. We exploit this property to develop a cross-validation scheme for selecting neighborhood size and trimming parameters, together with oracle inequalities quantifying the performance of the selected tuning parameters (Stone, 1974; Arlot and Celisse, 2010; Wainwright, 2019).
Importantly, the proposed modification does not compromise estimation accuracy. Under standard piecewise Lipschitz assumptions on the graphon, the LOO neighborhood smoother attains the same row-wise mean-squared error rates as classical neighborhood smoothing (Gao et al., 2015; Zhang et al., 2017). The additional computational cost of zeroing out a single row and column is negligible relative to the overall complexity of neighborhood construction.
Leave-one-out is a familiar device in regression, but its role here is not merely mechanical. Neighborhood selection depends on empirical two-hop geometry, which is itself random and sensitive to perturbations of the adjacency matrix. A naive leave-one-out scheme could destroy neighborhood locality and inflate bias. Our contribution is to show that zeroing a single row and column preserves neighborhood geometry with high probability: the trimmed global threshold and the $h$-nearest neighbor sets are stable under the LOO perturbation. This stability is the key technical ingredient that allows independence to be restored without sacrificing bias control.
@@
The remainder of the paper is organized as follows. Section 2 introduces the model, the Zhang-Levina-Zhu distance, and the leave-one-out neighborhood construction. Section 3 develops variance-adaptive concentration bounds, Berry-Esseen normal approximations, and confidence intervals for individual edge probabilities. Section 4 presents the honest cross-validation scheme and establishes an oracle inequality. Section 5 reports simulation results illustrating finite-sample behavior, and Section 6 concludes with a discussion of extensions and open problems.
@@ 2.1 LOO neighborhoods and estimator
Fix a neighborhood size $h=h_{n}$ (in theory $h\asymp\sqrt{n\log n}$) and a trimming quantile $\tau\in(0,1)$. For each row $i$, let $d_{i}^{*}$ denote the $h$-th order statistic of $\{d_{ij}:j\neq i\}$, and define the trimmed‑global threshold
$\bar{d}_{\tau}:=\mathrm{Quantile}_{\tau}(d_{1}^{*},\dots,d_{n}^{*}),$ (2.3)
the $\tau$-quantile of $\{d_{i}^{*}\}_{i=1}^{n}$. This construction yields an adaptive neighborhood: although the target size is fixed at $h$, the membership of the neighborhood is determined data‑dependently through the empirical ZLZ distances and the global threshold $\bar{d}_{\tau}$.
The preliminary neighborhood of node $i$ is
$\widetilde{\mathcal{N}}_{i}:=\{j\neq i:d_{ij}\leq\bar{d}_{\tau}\}.$
If $|\widetilde{\mathcal{N}}_{i}|\geq h$ we truncate to the $h$ closest indices by $d_{ij}$; if $|\widetilde{\mathcal{N}}_{i}|<h$ we augment by adding nearest neighbors until $|\mathcal{N}_{i}|=h$. Thus $|\mathcal{N}_{i}|=h$ for all $i$, but the membership of $\mathcal{N}_{i}$ is fully data‑adaptive.
The classical neighborhood smoother of Zhang et al. (2017) uses $\mathcal{N}_{i}$ computed from the full adjacency matrix $A$ to estimate the entire $i$th row of $P$. Because $\mathcal{N}_{i}$ depends on all columns of $A$, including column $j$, the weights used to average $\{A_{kj}:k\in\mathcal{N}_{i}\}$ depend on the very summands being averaged. Hence the centered terms $A_{kj}-P_{kj}$ are not independent conditional on the neighborhoods, preventing the direct application of variance‑adaptive concentration or Berry–Esseen theory.
To restore conditional independence when estimating $P_{ij}$, we adopt a leave‑one‑out construction. For fixed $(i,j)$ with $i\neq j$, let $A^{(-j)}$ denote the adjacency matrix obtained from $A$ by setting the $j$th row and column to zero. Recompute the distances in (2.2) using $A^{(-j)}$, obtaining $d_{i\ell}^{(-j)}$ for $\ell\neq i,j$. Using the same parameters $(h,\tau)$ and the global threshold $\bar{d}_{\tau}$, define
$\widetilde{\mathcal{N}}_{i}^{(-j)}:=\{\ell\neq i:d_{i\ell}^{(-j)}\leq\bar{d}_{\tau}\},$
and truncate or augment as before so that $|\mathcal{N}_{i}^{(-j)}|=h$. By constructing $\mathcal{N}_{i}^{(-j)}$ from $A^{(-j)}$, the random choice of neighbors for row $i$ does not depend on the edges $\{A_{kj}:k\in\mathcal{N}_{i}^{(-j)}\}$ used to estimate $P_{ij}$. This decoupling is the key structural device that restores conditional independence.
@@
By construction, $A^{(-j)}$ is obtained from $A$ by setting the $j$th row and column to zero. Hence $A^{(-j)}$ is a measurable function of $\{A_{k\ell}:k\neq j,\ell\neq j\}$, and so is $\mathcal{N}_{i}^{(-j)}$. Therefore, conditional on $P$, the random set $\mathcal{N}_{i}^{(-j)}$ is independent of the collection $\{A_{kj}:k\neq j\}$.
@@
3 Variance Bounds, Berry-Esseen Approximation, and Confidence Intervals
In this section we quantify the deviation of the LOO estimator $\tilde{P}_{ij}$ from the target $P_{ij}$. The analysis separates the estimation error into a variance component, arising from the randomness of observed edges given $P$, and a bias component, arising from approximating $P_{ij}$ by a local average over a neighborhood of $i$. We begin with a decomposition that will be used throughout the section. For $i\neq j$ write
@@ 3 Variance Bounds, Berry-Esseen Approximation, and Confidence Intervals
$U_{ij}$ $:=\frac{1}{h}\sum_{k\in\mathcal{N}_{i}^{(-j)}}(A_{kj}-P_{kj}),$ (3.1)
$B_{ij}$ $:=\frac{1}{h}\sum_{k\in\mathcal{N}_{i}^{(-j)}}(P_{kj}-P_{ij}).$ (3.2)
By Proposition 2.7, conditional on $\mathcal{N}_{i}^{(-j)}$ and $P$ the summands in $U_{ij}$ are independent, mean zero and bounded, so $U_{ij}$ behaves like an average of $h$ independent observations. The term $B_{ij}$ is a deterministic (given $P$ and $\mathcal{N}_{i}^{(-j)}$) smoothness bias that can be controlled using Assumption 2.1 and the neighborhood properties in Section 2.
3.1 Variance proxy and Bousquet inequality
We introduce a conditional variance proxy for $U_{ij}$:
$V_{ij}:=\mathrm{Var}\big(U_{ij}\mid\mathcal{N}_{i}^{(-j)},P\big)=\frac{1}{h^{2}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}(1-P_{kj}),$ (3.3)
so that $V_{ij}\leq 1/(4h)$. Note that $V_{ij}$ is measurable with respect to the $\sigma$-field generated by $(\mathcal{N}_{i}^{(-j)},P)$ and can be treated as fixed when conditioning on that field.
Lemma 3.1 (Bousquet inequality).
For any $t>0$,
$\mathbb{P}\Big(|U_{ij}|\geq\sqrt{2V_{ij}t}+\frac{t}{3h}\ \Big|\ \mathcal{N}_{i}^{(-j)},P\Big)\leq 2e^{-t}.$
Sketch of proof.
Conditional on $(\mathcal{N}_{i}^{(-j)},P)$ set $X_{k}=A_{kj}-P_{kj}$ for $k\in\mathcal{N}_{i}^{(-j)}$. The $X_{k}$ are independent, mean zero and bounded by 11. Let $S=\sum_{k\in\mathcal{N}_{i}^{(-j)}}X_{k}$, so that $U_{ij}=h^{-1}S$ and $\mathrm{Var}(S\mid\mathcal{N}_{i}^{(-j)},P)=h^{2}V_{ij}$. Apply Bousquet’s inequality for bounded independent summands (Bousquet, 2002; Vershynin, 2018) to $S$ and divide by $h$ to obtain the stated bound. ∎
Lemma 3.1 gives a variance‑adaptive, finite‑sample control of $U_{ij}$ in terms of the plug‑in quantity $V_{ij}$, which we estimate below.
@@ 3.2 Neighborhood bias
Lemma 3.2 (Neighborhood bias).
Let $A\in\{0,1\}^{n\times n}$ be the adjacency matrix and let $P$ be the corresponding probability matrix under the graphon model. For each index $j$ define the leave-one-out matrix $A^{(-j)}$ by setting the $j$th row and column to zero:
$A^{(-j)}_{k\ell}=\begin{cases}0,&\text{if }k=j\text{ or }\ell=j,\\ A_{k\ell},&\text{otherwise}.\end{cases}$
Let $\mathcal{N}_{i}^{(-j)}$ be the $h$-neighborhood of $i$ constructed from $M^{(-j)}=(A^{(-j)})^{2}/n$ using the trimmed ZLZ procedure with global threshold $\bar{d}_{\tau}$. Define
$B_{ij}=\frac{1}{h}\sum_{k\in\mathcal{N}_{i}^{(-j)}}(P_{kj}-P_{ij}).$
Under Assumptions 2.1 and 2.2 and with $h\asymp\sqrt{n\log n}$, there exists $C>0$ such that, with probability tending to one,
$\max_{1\leq i,j\leq n}|B_{ij}|\;\leq\;C\Big(\frac{\log n}{n}\Big)^{1/4}.$ (3.4)
Proof sketch.
A uniform concentration bound for the two-hop matrix follows from Bernstein’s inequality applied to the entries of $A^{2}/n$. In particular,
$\max_{i,k,\ell}\big|(M_{i\ell}-M_{k\ell})-\mathbb{E}[M_{i\ell}-M_{k\ell}]\big|\lesssim\sqrt{\frac{\log n}{n}}$
with probability 1o(1)1-o(1). Under Assumption 2.1, the population two-hop distance $\Delta_{ik}$ is Lipschitz in $|\xi_{i}-\xi_{k}|$, so nodes that are close in latent position have small population distance. Combining this latent geometry with the above concentration bound and the trimmed ZLZ construction shows that the empirical neighborhoods satisfy
$\max_{k\in\mathcal{N}_{i}}|\xi_{k}-\xi_{i}|\lesssim(\log n/n)^{1/4}$
with high probability. Replacing $A$ by $A^{(-j)}$ perturbs each entry of $M$ by at most $O(1/n)$, since only paths passing through $j$ are affected. Lemma B.2 ensures that such a perturbation changes the global threshold $\bar{d}_{\tau}$ and the resulting $h$-nearest neighbor sets in only a negligible number of positions. Consequently, the same latent proximity bound holds for the leave-one-out neighborhoods $\mathcal{N}_{i}^{(-j)}$. Because $h$ is Lipschitz in its first argument, the bound on $|\xi_{k}-\xi_{i}|$ implies
$|P_{kj}-P_{ij}|\lesssim(\log n/n)^{1/4}\qquad\text{for all }k\in\mathcal{N}_{i}^{(-j)}.$
Averaging over $h$ such terms yields the desired bias bound. ∎
3.3 Row‑wise risk
Combining the variance and bias controls yields a row‑wise risk bound.
Theorem 3.3 (Row‑wise risk rate).

### v3

@@ Abstract
Estimating latent edge probabilities in network models often relies on aggregating information from structurally similar nodes. However, using the same adjacency matrix to both determine node similarities and estimate probabilities induces complex statistical dependencies. This coupling invalidates standard inferential tools, such as concentration inequalities and central limit theorems, leaving a critical gap in entrywise uncertainty quantification for networks. To resolve this, we introduce a leave-one-out (LOO) decoupling strategy. When predicting a target edge, the local neighborhood is constructed from a reduced adjacency matrix with the target row and column entirely removed. This structural modification-rather than the specific smoothing mechanism-is the core innovation: it restores exact conditional independence between the data-driven neighborhood selection and the observed edges being aggregated. Leveraging this restored independence, we derive variance-adaptive empirical Bernstein bounds and establish uniform Berry-Esseen normal approximations with explicit rates. These results enable the construction of valid finite-sample and asymptotic confidence intervals for individual edge probabilities. Furthermore, the LOO architecture naturally admits an honest cross-validation scheme for tuning parameter selection, providing a complete and rigorous inferential framework while preserving strong mean-squared error performance.
@@ 1 Introduction
Graphon models have emerged as the canonical nonparametric framework for analyzing complex network data, providing a flexible representation of latent edge probabilities through a measurable kernel on the unit square (Borgs et al., 2007; Lovász, 2012). Over the past decade, a rich literature has developed around estimating the underlying probability matrix $P$ or the graphon $h$ from a single observed adjacency matrix $A$. Early foundational work focused heavily on the Stochastic Block Model (SBM) and its variants, utilizing modularity maximization, maximum likelihood, and spectral clustering to recover block probabilities (Bickel and Chen, 2009; Abbe, 2017; Rohe et al., 2011). As the field moved toward continuous, nonparametric graphons, modern estimators-including universal singular value decomposition (Chatterjee, 2015), sorting-based smoothing (Chan and Airoldi, 2014), and block-approximations (Gao et al., 2015)-have been proven to reliably achieve minimax-optimal rates for point estimation.
Among these, the neighborhood smoothing method introduced by Zhang et al. (2017) stands out. By adaptively aggregating structurally similar nodes based on empirical two-hop distances, neighborhood smoothing achieves optimal mean-squared error (MSE) without requiring the strict structural assumptions of block models or the selection of a global low-rank embedding.
However, a fundamental gap remains: the transition from optimal point estimation to rigorous statistical inference. Given an entrywise estimate $\widehat{P}_{ij}$, it is critically important for downstream scientific tasks to attach a valid measure of uncertainty to this prediction. Yet, existing inferential theory for networks has focused almost exclusively on low-dimensional or global functionals. Significant progress has been made in proving limit theorems for spectral embeddings (Sussman et al., 2012; Athreya et al., 2016), asymptotic normality for community parameters in SBMs (Bickel et al., 2013), and global goodness-of-fit testing (Lei, 2016). Conversely, entrywise uncertainty quantification-specifically, constructing valid confidence intervals and proving central limit theorems for individual edge probabilities under continuous graphons-has remained theoretically elusive. Recent attempts to bridge this gap have relied on computationally intensive resampling methods, such as network bootstrapping (Levin and Levina, 2025; Green and Shalizi, 2022), or generalized fiducial inference (Su et al., 2022), but exact distribution-free analytical bounds for neighborhood estimators are lacking.
The obstacle preventing entrywise inference is not merely technical, but structural. To achieve optimal MSE, neighborhood smoothing estimators (Zhang et al., 2017) rely on a data-driven aggregation of similar nodes. This creates a severe “double-dipping” problem: the procedure reuses the adjacency matrix twice. The matrix $A$ is first used to compute pairwise topological distances to select a neighborhood $\mathcal{N}_{i}$, and then the very same matrix $A$ is used to average the edges within that neighborhood. Consequently, the random neighborhood weights depend directly on the edges being averaged. Even after conditioning on the latent probability matrix $P$, the summands in the estimator are strongly coupled. This intractable dependence invalidates the foundational assumptions of classical concentration inequalities and Berry-Esseen normal approximations.
In this paper, we demonstrate that this probabilistic bottleneck can be cleanly bypassed via a simple but highly principled structural modification: leave-one-out (LOO) neighborhood smoothing.
When estimating a single entry $P_{ij}$, we perform a topological surgery on the graph by constructing the neighborhood of node $i$ using a reduced adjacency matrix $A^{(-j)}$ in which the $j$-th row and column have been entirely deleted. All pairwise distances and nearest-neighbor selections are computed on this reduced matrix. Only after the neighborhood $\mathcal{N}_{i}^{(-j)}$ is fixed do we average the original target edges $\{A_{kj}:k\in\mathcal{N}_{i}^{(-j)}\}$.
This decoupling is the central theoretical lever of the paper. By strictly separating the data used for neighborhood selection from the data used for estimation, we restore exact conditional independence among the centered summands. This topological decoupling bridges the gap between network estimation and classical empirical process theory. Specifically, under standard piecewise Lipschitz assumptions on the graphon, we leverage this restored independence to establish three main contributions:
  1. 1.
    Variance-Adaptive Concentration: We derive tight, finite-sample concentration inequalities based on empirical Bernstein bounds that automatically adapt to the local variance of the edge probabilities.
  2. 2.
    Asymptotic Normality: We establish uniform Berry-Esseen bounds with explicit convergence rates for the normalized estimation error, enabling the construction of valid normal-approximation confidence intervals for individual edge probabilities.
  3. 3.
    Honest Cross-Validation: The LOO structure naturally supports an unbiased cross-validation scheme for tuning the neighborhood size $h_{n}$.inequality.
Crucially, these inferential guarantees do not come at the cost of statistical efficiency. We prove that the proposed LOO estimator preserves the optimal row-wise mean-squared error rates of classical neighborhood smoothing.
@@
The remainder of the paper is organized as follows. Section 2 formalizes the graphon model, the Zhang-Levina-Zhu (ZLZ), Zhang et al. (2017) empirical distance, and our leave-one-out neighborhood construction. Section 3.2 establishes the foundational conditional independence, proving the variance-adaptive concentration bounds, Berry-Esseen normal approximations, and the validity of the resulting confidence intervals. Section 4 details the honest cross-validation scheme for tuning parameter selection. Section 5 reports comprehensive simulation results illustrating the finite-sample coverage and MSE performance, and Section 6 concludes with a discussion of extensions to directed, weighted, and degree-corrected settings.
@@ 2.1 LOO neighborhoods and estimator
For estimation of a fixed entry $(i,j)$, remove node $j$ and form the reduced adjacency matrix $A^{(-j)}$ obtained by deleting the $j$th row and column. Define the two-hop matrix
$M^{(-j)}=\frac{1}{n-1}A^{(-j)}A^{(-j)}.$ (2.3)
For node $i$ and node $k\neq i,j$, define the ZLZ distance as follows:
[…]
Let
$d_{i,(1)}^{(-j)}\leq\cdots\leq d_{i,(n-2)}^{(-j)},$
denote the order statistics of all nodes except $j$ to node $i$, namely $\{d_{ik}^{(-j)}:k\neq i,j\}$. Define the leave-one-out neighborhood of node $i$ by
$\mathcal{N}_{i}^{(-j)}=\left\{k\neq i,j:d_{ik}^{(-j)}\leq d_{i,(h_{n})}^{(-j)}\right\},$ (2.5)
that is, the $h_{n}$ nearest neighbors of $i$ according to the ZLZ two-hop distance computed on the reduced graph.
This construction preserves the original ZLZ neighborhood rule, with the sole modification that node $j$ is removed prior to distance computation.
The classical neighborhood smoother of Zhang et al. (2017) uses $\mathcal{N}_{i}$ computed from the full adjacency matrix $A$ to estimate the entire $i$th row of $P$. Because $\mathcal{N}_{i}$ depends on all columns of $A$, including column $j$, the weights used to average $\{A_{kj}:k\in\mathcal{N}_{i}\}$ depend on the very summands being averaged. Hence the centered terms $A_{kj}-P_{kj}$ are not independent conditional on the neighborhoods, preventing the direct application of variance‑adaptive concentration or Berry-Esseen theory.
@@
By construction, $A^{(-j)}$ is obtained from $A$ by removing the $j$th row and column. Hence $A^{(-j)}$ is a measurable function of $\{A_{k\ell}:k\neq j,\ell\neq j\}$, and so is $\mathcal{N}_{i}^{(-j)}$. Therefore, conditional on $P$, the random set $\mathcal{N}_{i}^{(-j)}$ is independent of the collection $\{A_{kj}:k\neq j\}$.
@@
3 Consistency
In this section we quantify the deviation of the LOO estimator $\tilde{P}_{ij}$ from the target $P_{ij}$. The analysis separates the estimation error into a variance component, arising from the randomness of observed edges given $P$, and a bias component, arising from approximating $P_{ij}$ by a local average over a neighborhood of $i$. We begin with a decomposition that will be used throughout the section. For $i\neq j$ write For $i\neq j$, write
@@ 3 Consistency
$U_{ij}^{(-j)}$ $:=\tilde{P}_{ij}-\frac{1}{h}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}=\frac{1}{h}\sum_{k\in\mathcal{N}_{i}^{(-j)}}(A_{kj}-P_{kj}),$ (3.1)
$B_{ij}^{(-j)}$ $:=\frac{1}{h}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}-P_{ij}.$ (3.2)
By Proposition 2.7, conditional on $\mathcal{N}_{i}^{(-j)}$ and $P$ the summands in $U_{ij}^{(-j)}$ are independent, mean zero and bounded, so $U_{ij}^{(-j)}$ behaves like an average of $h$ independent observations. The term $B_{ij}^{(-j)}$ is a deterministic (given $P$ and $\mathcal{N}_{i}^{(-j)}$) smoothness bias that can be controlled using Assumption 2.1 and the neighborhood properties in Section 2.
@@ 3.1 Neighborhood bias
Lemma 3.1 (Uniform two-hop stability under deletion of one vertex).
Let $A\in\{0,1\}^{n\times n}$ be the adjacency matrix of an undirected simple graph with $A_{ii}=0$. For each $j\in[n]$, let $A^{(-j)}$ be the $(n-1)\times(n-1)$ adjacency matrix obtained by deleting the $j$th row and column (i.e., removing vertex $j$). Define the normalized two-hop matrices
$M:=\frac{1}{n}A^{2},\qquad M^{(-j)}:=\frac{1}{n-1}\big(A^{(-j)}\big)^{2}.$
For each $j$, embed $M^{(-j)}$ into an $n\times n$ matrix by placing it on the index set $n]\setminus\{j\}$ and setting entries involving index $j$ to zero. Then, deterministically,
$\sup_{1\leq j\leq n}\ \|M-M^{(-j)}\|_{\max}\ \leq\ \frac{3}{n},$
where $\|B\|_{\max}:=\max_{1\leq i,\ell\leq n}|B_{i\ell}|$ denotes the entrywise max norm. In particular, for every $j\in[n]$ and all $i,\ell\in[n]\setminus\{j\}$,
$\big|M_{i\ell}-M^{(-j)}_{i\ell}\big|\leq\frac{3}{n}.$
Proof.
Fix $j\in[n]$ and $i,\ell\in[n]\setminus\{j\}$. Expanding the $(i,\ell)$ entry gives
$(A^{2})_{i\ell}=\sum_{m=1}^{n}A_{im}A_{m\ell},\qquad\big(A^{(-j)}\big)^{2}_{i\ell}=\sum_{m\neq j}A_{im}A_{m\ell},$
hence
$(A^{2})_{i\ell}-\big(A^{(-j)}\big)^{2}_{i\ell}=A_{ij}A_{j\ell},$
so $\big|(A^{2})_{i\ell}-\big(A^{(-j)}\big)^{2}_{i\ell}\big|\leq 1$. Therefore,
[…]
$\leq\left|\frac{1}{n}(A^{2})_{i\ell}-\frac{1}{n}\big(A^{(-j)}\big)^{2}_{i\ell}\right|+\left|\frac{1}{n}-\frac{1}{n-1}\right|\big(A^{(-j)}\big)^{2}_{i\ell}$
$\leq\frac{1}{n}+\frac{1}{n(n-1)}\big(A^{(-j)}\big)^{2}_{i\ell}.$
Since $\big(A^{(-j)}\big)^{2}_{i\ell}=\sum_{m\neq j}A_{im}A_{m\ell}\leq n-1$, we get
$\big|M_{i\ell}-M^{(-j)}_{i\ell}\big|\leq\frac{1}{n}+\frac{n-1}{n(n-1)}=\frac{2}{n}.$
Thus the bound holds for every $j$ (in fact with constant $2$), and taking the supremum over $j$ yields $\|M-M^{(-j)}\|_{\max,[n]\setminus\{j\}}:=\max_{i,\ell\in[n]\setminus\{j\}}|M_{i\ell}-M^{(-j)}_{i\ell}|\leq 2/n$. ∎
Lemma 3.2 (Two-hop concentration is preserved under leave-one-out deletion).
Let $A\in\{0,1\}^{n\times n}$ be generated from the graphon model (2.1) with probability matrix $P$. Define
$M:=\frac{1}{n}A^{2},\qquad M^{\star}:=\mathbb{E}\!\left[M\,\middle|\,P\right],$
and for each $j\in[n]$ let $A^{(-j)}$ be the adjacency matrix obtained by deleting the $j$th row and column. Define the leave-one-out two-hop matrices
$M^{(-j)}:=\frac{1}{n-1}\big(A^{(-j)}\big)^{2},\qquad M^{(-j)\star}:=\mathbb{E}\!\left[M^{(-j)}\,\middle|\,P\right].$
Embed $M^{(-j)}$ and $M^{(-j)\star}$ into $n\times n$ matrices on the index set $n]\setminus\{j\}$ (and set entries involving $j$ to zero).
Then, for every $n$,
$\sup_{1\leq j\leq n}\ \big\|M^{(-j)}-M^{(-j)\star}\big\|_{\max}\ \leq\ \|M-M^{\star}\|_{\max}\ +\ \frac{C}{n},$ (3.3)
for a universal constant $C>0$. In particular, if
$\|M-M^{\star}\|_{\max}=\mathcal{O}_{\mathbb{P}}\!\Big(\sqrt{\tfrac{\log n}{n}}\Big),$
(as established for the ZLZ two-hop matrix in Zhang et al. (2017) and its supplementary material), then
$\sup_{1\leq j\leq n}\ \big\|M^{(-j)}-M^{(-j)\star}\big\|_{\max}=\mathcal{O}_{\mathbb{P}}\!\Big(\sqrt{\tfrac{\log n}{n}}\Big).$
Proof.
Fix $j\in[n]$. By the triangle inequality,
$\big\|M^{(-j)}-M^{(-j)\star}\big\|_{\max}$ $\leq\big\|M^{(-j)}-M\big\|_{\max}+\big\|M-M^{\star}\big\|_{\max}+\big\|M^{\star}-M^{(-j)\star}\big\|_{\max}.$
The first term is deterministic and satisfies
$\sup_{1\leq j\leq n}\|M^{(-j)}-M\|_{\max}\leq\frac{C_{1}}{n}$
by Lemma 3.1. For the last term, note that
$M^{\star}=\mathbb{E}[M\mid P],\qquad M^{(-j)\star}=\mathbb{E}[M^{(-j)}\mid P],$
so Jensen’s inequality and Lemma 3.1 give
$\sup_{1\leq j\leq n}\|M^{\star}-M^{(-j)\star}\|_{\max}=\sup_{1\leq j\leq n}\big\|\mathbb{E}[M-M^{(-j)}\mid P]\big\|_{\max}\leq\sup_{1\leq j\leq n}\mathbb{E}\big[\|M-M^{(-j)}\|_{\max}\mid P\big]\leq\frac{C_{1}}{n}.$
Combining the bounds yields (3.3) with $C=2C_{1}$. Finally, since $n^{-1}=o\!\big(\sqrt{(\log n)/n}\big)$, the stated rate transfer follows immediately. ∎
Proposition 3.3 (ZLZ neighborhood smoothing: row-risk (black box)).
Work under the assumptions of Zhang, Levina, and Zhu (2017, Biometrika) and choose the bandwidth $h\asymp\sqrt{n\log n}$ as in their Theorem 1. Let $N_{i}$ be the ZLZ neighborhood of node $i$ constructed from the two-hop matrix $M=A^{2}/n$ via their distance and thresholding rule, and define the ZLZ estimator
$\tilde{P}^{\mathrm{ZLZ}}_{ij}:=\frac{1}{|N_{i}|}\sum_{k\in N_{i}}A_{kj}.$